# Bridged Clustering for Representation Learning: Semi-Supervised Sparse Bridging

### v1

@@ Abstract
We introduce Bridged Clustering, a semi-supervised framework to learn predictors from any unpaired input $\mathcal{X}$ and output $\mathcal{Y}$ dataset. Our method first clusters $\mathcal{X}$ and $\mathcal{Y}$ independently, then learns a sparse, interpretable bridge between clusters using only a few paired examples. At inference, a new input $x$ is assigned to its nearest input cluster, and the centroid of the linked output cluster is returned as the prediction $\hat{y}$. Unlike traditional SSL, Bridged Clustering explicitly leverages output-only data, and unlike dense transport-based methods, it maintains a sparse and interpretable alignment. Through theoretical analysis, we show that with bounded mis-clustering and mis-bridging rates, our algorithm becomes an effective and efficient predictor. Empirically, our method is competitive with SOTA methods while remaining simple, model-agnostic, and highly label-efficient in low-supervision settings.
@@ 1 Introduction
Analyzing modern datasets often entails reconciling rich but unaligned measurements collected for different purposes. For example, researchers may inherit a large morphological collection of museum specimen records alongside an independent genetic sequence repository such as GenBank, without reliably coordinated specimen identifiers to join them (Mulcahy, 2022). This scenario is very common across domains such as biomedicine, ecology, and the social sciences, where modern data ecosystems rarely provide perfectly paired records for every variable of interest (Putrama & Martinek, 2024; Wafa’Za’al Alma’aitah et al., 2024). Building supervised models in these settings is challenging, often requiring expensive relabeling or non-scalable heuristic alignment strategies.
Semi-supervised learning (SSL) provides a natural approach when labeled pairs are scarce. By leveraging large sets of unlabeled inputs, SSL methods such as graph-based label propagation (Zhu & Ghahramanirh, 2002), consistency regularization techniques like Mean Teacher (Tarvainen & Valpola, 2017) and Temporal Ensembling (Laine & Aila, 2016), and hybrid approaches such as MixMatch (Berthelot et al., 2019) and FixMatch (Sohn et al., 2020) can dramatically reduce annotation costs. Similarly, transductive regressors like LapRLS (Belkin et al., 2006) have proven effective in settings with limited supervision. However, these approaches assume that each unlabeled example is an input $x$ whose label $y$ is merely hidden (Van Engelen & Hoos, 2020; Yang et al., 2022). This assumption breaks in the presence of output-only data, which is common in practice. Output-only data includes, for example, unpaired DNA barcodes collected independently of morphological images (Ratnasingham et al., 2024), clinical outcome records disconnected from patient biosignals (Sayers et al., 2019), satellite imagery without matching weather station logs, audio recordings lacking textual transcriptions, or microscopy images without corresponding gene-sequencing profiles. Thus, to fully capitalize on available data, a predictive model must learn not only from unlabeled inputs but also from the structure and abundance of outputs. Motivated by this challenge, we ask:
@@
Several related frameworks have been proposed to connect disjoint input and output spaces. Transport-based alignment methods leverage output-only data via distributional coupling: entropic optimal transport with barycentric mapping (EOT-BM) uses a cross-domain cost $c(x,y)$ to couple $\mathcal{X}$ and $\mathcal{Y}$ and predicts by barycentric projection, while Gromov–Wasserstein (GW) aligns intra-space geometries when no cross-domain metric is available (Cuturi, 2013; Mémoli, 2011; Peyré et al., 2019). Unmatched regression and data fusion tackle separate samples of $X$ and $Y$ under structural assumptions (e.g., linear or monotone links, kernel embeddings) (Azadkia & Balabdaoui, 2024b; Evans et al., 2021; Robbins et al., 2024; Hsu & Ramos, 2019; Chau et al., 2021). Generative alternatives such as mixture-of-regressions via EM (Dempster et al., 1977) alternate imputations of missing inputs or outputs. Multi-view learning and co-training align modalities but require paired views per example (Blum & Mitchell, 1998; Bickel & Scheffer, 2004; Rey & Roth, 2012b). Our approach differs by being semi-supervised at the cluster level: we independently structure each space using unpaired data, then resolve the input–output association with a small paired set, yielding a sparse, interpretable bridge and linear-time scaling in $n$ once $k\ll n$.
We introduce Bridged Clustering, a simple and general SSL algorithm that builds a predictor from unpaired datasets using only a small set of matched examples (Algorithm 1). In a nutshell, it:
  1. 1.
    Independently cluster the input and output spaces.
  2. 2.
    Learn a cluster-level bridge from the small paired set.
  3. 3.
    Predict by assigning a new input to its nearest input cluster and returning the centroid of the linked output cluster.
Bridged Clustering is model-agnostic, as any embedding model (e.g., ResNet, BERT) and standard clustering algorithm (e.g., $k$-means, spectral) can be used, and requires as few as one paired example per active input cluster. By leveraging structure in both modalities and bridging them through a minimal set of task-specific labeled pairs, Bridged Clustering enables accurate prediction without the need for large-scale supervision. Moreover, since both unlabeled $\mathcal{X}$ and unlabeled $\mathcal{Y}$ contribute to refining the cluster centroids, the estimator’s risk decreases with the total pool size, even in regimes with extremely limited pairing. We demonstrate the effectiveness of this method across diverse domains: vision, language, and bioinformatics.
Design simplicity. Our method is minimal, using standard clustering in both spaces and a sparse majority-vote bridge learned from a handful of labeled pairs. It is model-agnostic at both the encoder and clustering algorithm levels, and interpretable through cluster-to-cluster correspondences.
Positioning. Like transport-based alignments and other unmatched regression methods, our method leverages output-only data; unlike them, we use a tiny paired set to learn a sparse, cluster-to-cluster bridge with explicit error controls. Our contributions are:
  • Robust learning under sparse supervision by using large pools of unlabeled inputs and outputs to refine cluster centroids and training the bridge with only a few paired examples.
  • Leveraging independently collected heterogeneous data by unifying disjoint input and output sources, so that structure in one modality informs prediction in the other.
  • Bidirectional inference across domains via a cluster-level mapping that enables prediction of outputs from inputs and inputs from outputs without additional paired data.
  • Simplicity, interpretability, and efficiency. The pipeline uses only standard clustering and a majority-vote bridge, yielding a sparse, interpretable cluster map and centroid predictors. This design is model-agnostic (plug-in encoders and clustering subroutines) and leads to linear-time scaling once $C\!\ll\!n$, as formalized in our runtime analysis (Section 5.3).
Our findings suggest a new research axis for SSL: jointly exploiting input-only and output-only data by independent clustering. Bridged Clustering opens the door to leveraging the many unpaired available repositories without incurring prohibitive annotation costs. [^1] We release code and data: https://github.com/patrickpxye/bc.
@@ 2 Related Work
Transductive and Inductive SSL:
Early graph-based methods augment labeled data with unlabeled inputs by diffusing labels over a similarity graph on $\mathcal{X}$ (Zhou et al., 2003); Examples include Gaussian random-walk propagation and Laplacian Regularized Least Squares (LapRLS; Belkin et al. (2006)). Modern consistency-regularization methods like Temporal Ensembling (Laine & Aila, 2016) and the Mean Teacher model (Tarvainen & Valpola, 2017) improve predictions by enforcing stability under input perturbations. MixMatch (Berthelot et al., 2019) unified label guessing with MixUp to encourage low-entropy regions, with FixMatch (Sohn et al., 2020) simplifying this via weak/strong augmentations with confidence-thresholded pseudo-labels.
Dedicated SSL methods for regression remain rare. Twin Neural Network Regression (TNNR) (Wetzel et al., 2022) uses pairwise difference targets and unpaired anchor loops for self-consistency. Recent advances include Uncertainty-Consistent Variational Model Ensembling UCVME (Dai et al., 2023), which models aleatoric noise for calibrated pseudo-labels, and RankUp (Huang et al., 2024), framing regression as an auxiliary ranking task to leverage classification SSL techniques.
Graph neural networks (GNNs) provide additional transductive baselines. Graph Convolutional Networks (GCN; Kipf & Welling (2016)) combine neighborhood aggregation with partial supervision, while Adaptive Diffusion Convolution (ADC; Zhao et al. (2021)) learns hop-wise diffusion weights to mitigate over-smoothing. Unlike Bridged Clustering, these methods only use unlabeled inputs and cannot leverage external output-only data.
Multi-View and Dependency-Seeking Clustering:
Co-training trains two classifiers on disjoint feature views, exchanging pseudo-labels to leverage unlabeled data given conditional independence (Blum & Mitchell, 1998). Later works explore PAC guarantees (Dasgupta & Monteleoni, 2005), relaxed expansions (Balcan et al., 2004), and practical enhancements (Nigam et al., 2000).
Dependency-Seeking Clustering (DSC) groups multi-view data by maximizing inter-view dependence (Rey & Roth, 2012a). Bayesian mixtures of local CCAs (Viinikanoja et al., 2010) and Canonical Least-Squares extend DSC by partitioning data via mixtures of linear CCA sub-models (Gitman et al., 2017). Unlike Bridged Clustering, all require paired views and cannot use disjoint $\mathcal{X}$- or $\mathcal{Y}$-only corpora.
Regression with Unmatched or Permuted Samples:
Recent statistical work on unmatched regression addresses separate, unordered datasets of $\mathcal{X}$ and $\mathcal{Y}$. Azadkia & Balabdaoui (2024a) cast linear regression as a deconvolution problem, establishing consistency and parametric rates. Others handle monotone or isotonic structures under unknown permutations (Carpentier & Schlüter, 2016; Rigollet & Weed, 2019). Bayesian data fusion approaches use kernel mean embeddings (Hsu & Ramos, 2019; Chau et al., 2021). Mixture-of-Regressions is trained via Expectation-Maximization (Dempster et al., 1977), which jointly imputes missing $\mathcal{X}$ or $\mathcal{Y}$ through iterations. Although this EM approach unifies unpaired and paired data, it does not leverage cluster geometry to sharpen predictions or produce explicit sparse cluster-level associations, setting Bridged Clustering apart.
@@
Optimal transport (OT) aligns empirical measures across domains using a cross-space cost $c(x,y)$ and can produce predictions via barycentric mapping of the learned coupling; entropic regularization enables fast Sinkhorn solvers (Cuturi, 2013; Peyré et al., 2019). When no cross-domain metric is available, Gromov–Wasserstein (GW) matches intra-space distances to recover a soft correspondence between metric-measure spaces (Mémoli, 2011; Peyré et al., 2016). Both families exploit output-only data through distributional structure. Our approach differs in using a small paired set to resolve mode identities and learning an explicit, sparse cluster bridge that scales linearly once $k\!\ll\!n$, with guarantees tailored to clustering-based estimation rather than global transport objectives. Whereas OT/GW learn dense, global couplings, our bridge is explicit and sparse, trading global transport for cluster-level interpretability and linear-time inference once $C\!\ll\!n$.
@@
In NLP, unsupervised machine translation aligns monolingual embeddings and relies on back-translation to bridge languages (Lample et al., 2017). In computer vision, CycleGAN and successors learn bidirectional mappings between image domains via cycle-consistency constraints (Zhu et al., 2017; Huang et al., 2018). Though operating without paired instances, these systems train end-to-end neural generators and require reconstructive losses in both directions, while Bridged Clustering is model-agnostic and learns only a sparse cluster-level association matrix using few labeled pairs.
3 Problem Formulation
We formalize Bridged Clustering’s prediction task and identify conditions under which it achieves bounded excess risk, even with limited paired data. Without loss of generality, we focus on predicting an output $Y$ from a given input $X$; the reverse direction (predicting $X$ from $Y$) is symmetric and follows by swapping roles. We now introduce key assumptions and notations.
Consider a population of samples partitioned into $C$ latent groups. While the criteria defining these groups are unknown, we assume the existence of $C$ distinct categories. Let $T\in[C]\!:=\!\{1,\dots,C\}$ be a latent categorical variable drawn with probabilities $(\pi_{1},\dots,\pi_{C})$. For each group $t$, let $P_{t}^{X}$ denote the distribution over input features and $P_{t}^{Y}$ the distribution over output features.
Conditional on $T=t$ we observe a pair: $(X,Y)\;\big|\;T=t\;\sim\;P_{t}^{X}\times P_{t}^{Y},$ and let
$\mu_{t}^{X}:=\mathbb{E}[X\mid T=t],\;\mu_{t}^{Y}:=\mathbb{E}[Y\mid T=t].$
@@ 3 Problem Formulation
  • a small paired sample $\mathcal{S}\!=\!\{(x^{\prime}_{j},y^{\prime}_{j})\}_{j=1}^{k}$, with $k\ll n_{X},n_{Y}$, drawn i.i.d. from joint $P=\sum_{t}\pi_{t}(P_{t}^{X}\!\times P_{t}^{Y})$.
@@ 4 Bridged Clustering Algorithm
sparse supervised set $\mathcal{S}=\{(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{k}$; Test-set $\mathcal{X_{\text{test}}}=\{x_{1},x_{2},\ldots,x_{|\mathcal{X_{\text{test}}}|}\}$
(Note: $\mathcal{X_{\text{test}}}=\mathcal{X}$ in transductive setting)
@@
1: (Optional) Enlarge $\mathcal{X}$ with the $x$-features of $\mathcal{S}$, enlarge $\mathcal{Y}$ with the $y$-features of $\mathcal{S}$.
2:Clustering in $\mathcal{X}$:
3: Apply a clustering algorithm to $\mathcal{X}$ to obtain cluster assignments $\mathcal{C}_{\mathcal{X}}$.
4:Clustering in $\mathcal{Y}$:
5: Apply a clustering algorithm to $\mathcal{Y}$ to obtain cluster assignments $\mathcal{C}_{\mathcal{Y}}$.
6:Bridge Learning:
7: Using the supervised set $\mathcal{S}$, learn a mapping $\hat{A}_{x\to y}$ between clusters in $\mathcal{X}$ and clusters in $\mathcal{Y}$.
8:Prediction:
9:for each sample $x_{i}$ in $\mathcal{X_{\text{test}}}$ do
10:  Assign $x_{i}$ to a cluster $c_{x}=\mathcal{C}_{\mathcal{X_{\text{test}}}}(x_{i})$.
11:  Find corresponding cluster in $\mathcal{Y}$: $c_{y}=\hat{A}_{x\to y}(c_{x})$.
12:  Predict $\hat{y}_{i}$ as the centroid of cluster $c_{y}$ in $\mathcal{Y}$.
13:end for
14:return Predicted outputs $\hat{\mathcal{Y}}$ for test inputs
@@
We first run a $k$-clustering algorithm to partition the dataset into $k$ groups separately on $\mathcal{X}$ and $\mathcal{Y}$ using $C$ centers: $\hat{c}_{X}:\mathbb{R}^{d}\!\to[C],\hat{c}_{Y}:\mathbb{R}^{d^{\prime}}\!\to[C].$ Because cluster labels are only defined up to a permutation of indices, we evaluate their quality via the permutation-invariant mis-clustering rates:
$\varepsilon_{X}:=\min_{\sigma:[C]\!\to\![C]}\Pr\bigl[\sigma(\hat{c}_{X}(X))\neq T\bigr],\varepsilon_{Y}:=\min_{\sigma:[C]\!\to\![C]}\Pr\bigl[\sigma(\hat{c}_{Y}(Y))\neq T\bigr].$
Let $\sigma_{X},\sigma_{Y}$ denote permutations that attain these minima.

### v3

# Bridged Clustering: Semi-Supervised Sparse Bridging

@@ Abstract
We introduce Bridged Clustering, a semi-supervised framework to learn predictors from any unpaired input $\mathcal{X}$ and output $\mathcal{Y}$ dataset. Our method first clusters $\mathcal{X}$ and $\mathcal{Y}$ independently, then learns a sparse, interpretable bridge between clusters using only a few paired examples. At inference, a new input $x$ is assigned to its nearest input cluster, and the centroid of the linked output cluster is returned as the prediction $\hat{y}$. Unlike traditional SSL, Bridged Clustering explicitly leverages output-only data, and unlike dense transport-based methods, it maintains a sparse and interpretable alignment. Through theoretical analysis, we show that with bounded mis-clustering and mis-bridging rates, our algorithm becomes an effective and efficient predictor. Empirically, our method is competitive with SOTA methods while remaining simple, model-agnostic, and label-efficient in low-supervision settings.
Machine Learning, Semi-supervised Learning, Clustering
Patrick Peixuan Ye Chen Shani Ellen Vitercik
Stanford University Stanford University Stanford University
pxye@stanford.edu cshani@stanford.edu vitercik@stanford.edu
[…]
Figure 1: Bridged Clustering overview: (1) Cluster $\mathcal{X}$ and $\mathcal{Y}$ independently, (2) learn a sparse cluster bridge $\hat{A}$ from $k$ paired samples by majority vote, (3) predict by mapping $x$ to its input cluster and returning the centroid of the bridged output cluster.
@@ 1 Introduction
Modern data analytics often involve reconciling rich but unaligned data collected for different purposes. For example, researchers may study a large morphological collection of museum specimen records alongside an independent genetic sequence repository such as GenBank (Sayers et al., 2019), without reliably coordinated specimen identifiers to join them (Mulcahy, 2022). This scenario is common across domains such as biomedicine, ecology, and the social sciences, where modern data ecosystems rarely provide perfectly paired records for every variable of interest (Putrama and Martinek, 2024; Wafa’Za’al Alma’aitah et al., 2024). Building supervised models in these settings often requires costly annotation, which becomes challenging at scale.
When labeled pairs are scarce, semi-supervised learning (SSL) is a natural approach, leveraging large pools of unlabeled inputs alongside limited supervised samples. Standard SSL assumes that each unlabeled input $x\in\mathcal{X}$ has a hidden label $y\in\mathcal{Y}$ (Van Engelen and Hoos, 2020; Yang et al., 2022), but this is often violated in multimodal settings, where datasets may include output-only data with no corresponding $x$. Examples for output-only data include: unpaired DNA barcodes collected independently of morphological images (Ratnasingham et al., 2024), clinical outcome records separate from patient biosignals (Mark, 2016), satellite imagery without matching weather station logs, audio recordings lacking textual transcriptions, or microscopy images without corresponding gene-sequencing profiles. In these settings, fully exploiting the available data requires learning not only from unlabeled inputs, but also from the distributional structure and abundance of unlabeled outputs. Motivated by this challenge, we ask:
@@
Key challenges.
Extending SSL to settings with unlinked inputs and outputs raises several challenges. First, SSL objectives rely on unlabeled inputs and instance-level pairing, and therefore cannot be directly applied when outputs are observed independently of inputs. Second, while output-only data encode valuable structural information about $\mathcal{Y}$, this information is not grounded in $\mathcal{X}$, making it unclear how it should influence prediction. Third, approaches that align inputs and outputs only at the marginal or distributional level risk coarse solutions. Finally, any practical method must exploit unlinked data without introducing substantial computational overhead or strong modeling assumptions.
Several lines of work partially address the challenges above. Transport-based alignment methods exploit output-only data by matching marginal distributions across spaces, either via cross-domain costs or by aligning intra-space geometries (Mémoli, 2011; Cuturi, 2013; Peyré and Cuturi, 2019). However, these approaches operate at the distributional level and may yield degenerate or overly coarse alignments. Related work on unmatched regression leverages separate samples of inputs and outputs under restrictive structural assumptions (Carpentier and Schlüter, 2016; Rigollet and Weed, 2019; Azadkia and Balabdaoui, 2024). Finally, generative and multi-view methods can effectively impute missing modalities, but typically require strong parametric assumptions or complete instance-level pairing (Dempster et al., 1977; Blum and Mitchell, 1998; Bickel and Scheffer, 2004).
We address these challenges by operating at the cluster level: we independently exploit unpaired data to structure $\mathcal{X}$ and $\mathcal{Y}$, and use a small paired set to learn a sparse, interpretable bridge between them.
Bridged Clustering is a model-agnostic SSL framework for learning from unpaired inputs and outputs with minimal paired supervision. It clusters each space independently, learns a sparse cluster-level association from a few paired examples, and predicts by mapping inputs to their linked output clusters; exploiting structure in both modalities reduces predictive risk as unpaired data increase. We evaluate Bridged Clustering across vision, language, and bioinformatics domains. Our contributions are:
  • We introduce a cluster-level SSL framework that learns from unpaired input-only and output-only data using minimal paired supervision.
  • We show how independently collected heterogeneous datasets can be unified through sparse cluster-level bridges, enabling effective prediction without instance-level alignment.
  • We present a simple, interpretable, and efficient algorithm that is model-agnostic, scales linearly in the number of examples, and yields transparent cluster-to-cluster correspondences.
Our results point to a complementary direction for SSL: jointly exploiting input-only and output-only data through independent structure discovery. [^1] We release code: https://github.com/patrickpxye/bc.
@@ 2 Related Work
Semi-supervised and Self-supervised Learning:
By leveraging large sets of unlabeled inputs, SSL methods such as graph-based label propagation (Zhu and Ghahramani, 2002; Zhou et al., 2003), transductive regressors like Laplacian Regularized Least Squares (LapRLS; Belkin et al. (2006)), consistency regularization techniques like Mean Teacher (Tarvainen and Valpola, 2017) and Temporal Ensembling (Laine and Aila, 2016), as well as hybrid approaches such as MixMatch (Berthelot et al., 2019) and FixMatch (Sohn et al., 2020) can all dramatically reduce annotation costs. For regression problems specifically, recent advances include Twin Neural Network Regression (TNNR) (Wetzel et al., 2022), Uncertainty-Consistent Variational Model Ensembling (UCVME) (Dai et al., 2023) and and RankUp (Huang et al., 2024). Unlike Bridged Clustering, these methods only use unlabeled inputs and do not leverage additional output-only data. Existing SSL objectives do not simultaneously consider unsupervised output alongside unlabeled inputs, even when abundant such data exist.
Self-supervised contrastive methods like SimCLR (Chen et al., 2020) involve unlabeled multi-view augmentations, and deep clustering methods such as SCAN (Van Gansbeke et al., 2020) and NNM (Dang et al., 2021) also rely on multi-view consistency. These approaches use multiple augmented views, but operate within a single input modality, while our model jointly targets inputs and outputs.
Multi-View Learning:
Multi-view and co-training methods explore the idea of learning from multiple distinct feature spaces, conceptually close to our setting. Related work include classical co-training and its PAC-style extensions (Blum and Mitchell, 1998; Dasgupta et al., 2001; Balcan et al., 2004; Balcan and Blum, 2010), dependency-seeking clustering approaches (Rey and Roth, 2012), mixture-based multi-view models such as Bayesian local CCA mixtures (Viinikanoja et al., 2010), and canonical least-squares clustering (Lei et al., 2017). While multi-view learning assumes co-occuring views $(X,Y)$ for each training instance, Bridged Clustering solves a different problem of learning a mapping from $X$ to $Y$ where most instances are unpaired. Unlike Bridged Clustering, these methods require paired views and therefore cannot use the disjoint $\mathcal{X}$- or $\mathcal{Y}$-only corpora in our problem setting.
Regression with Unmatched Samples:
Recent statistical work on unmatched regression address disjoint $\mathcal{X}$ and $\mathcal{Y}$ data. These work primarily rely on strong structural assumptions to recover input-output correspondence. Azadkia and Balabdaoui (2024) handle linear regression as a deconvolution problem, and other related work assume monotone or isotonic structures (Carpentier and Schlüter, 2016; Rigollet and Weed, 2019) to tackle unmatched regression. Another line of generative methods assumes latent parametric structure, such as mixture-of-regression optimized via Expectation–Maximization (EM) (Dempster et al., 1977). Bridged Clustering makes no such assumptions that these methods are tied to, while solving the same problem for aligning disjoint inputs and outputs.
@@
Optimal transport (OT) aligns input-output domains using a cross-space cost $c(x,y)$ and can produce predictions via barycentric mapping (Cuturi, 2013; Peyré and Cuturi, 2019). When cross-domain metric is unavailable, Gromov–Wasserstein (GW) matches intra-space distances to recover a soft correspondence between spaces (Mémoli, 2011; Peyré et al., 2016). Both methods exploit output-only data through distributional structure, prone to coarse alignments and computational burdens upon scaling (Peyré and Cuturi, 2019). Our approach differs in using a small paired set to learn a sparse cluster-bridge mapping. Whereas these methods learn dense, global couplings, our bridge is explicitly sparse, with cluster-level interpretability, and light computational overhead through only clustering and majority-voting (An asymptomatic efficiency comparison is in Appendix D).
@@
In NLP, unsupervised machine translation aligns monolingual embeddings and relies on back-translation to bridge languages (Lample et al., 2017). In computer vision, methods like CycleGAN learn bidirectional mappings between image domains via cycle-consistency constraints (Zhu et al., 2017; Huang et al., 2018). Although these systems operate without paired instances, they train end-to-end neural generators and typically rely on reconstructive losses in both directions, while Bridged Clustering is model-agnostic and learns a light cluster-level association using few labeled pairs.
3 Notation
In this section, we introduce our notation. The input features are denoted by random variable $X\in\mathbb{R}^{d}$ and the output features are denoted by random variable $Y\in\mathbb{R}^{d^{\prime}}.$
We assume that $X$ and $Y$ are mixtures of $C$ latent groups. While the criteria defining these groups are unknown, we assume the existence of $C$ distinct categories. Let
$T\in[C]\!:=\!\{1,\dots,C\}$
be a latent categorical variable drawn with probabilities $\pi_{1},\dots,\pi_{C}$. For each group $t$, let $P_{t}^{X}$ and $P_{t}^{Y}$ denote the distributions over $X$ and $Y$. Conditional on $T=t$, we observe a pair:
$(X,Y)\;\big|\;T=t\;\sim\;P_{t}^{X}\times P_{t}^{Y},$
and let $\mu_{t}^{X}:=\mathbb{E}[X\mid T=t]$ and $\mu_{t}^{Y}:=\mathbb{E}[Y\mid T=t].$
@@ 3 Notation
  • a paired sample $\mathcal{S}\!=\!\{(x^{\prime}_{j},y^{\prime}_{j})\}_{j=1}^{k}$, with $k\ll n_{X},n_{Y}$, drawn i.i.d. from joint $P=\sum_{t}\pi_{t}(P_{t}^{X}\!\times P_{t}^{Y})$.
The goal is to predict an output $Y$ from a given input $X$; the reverse direction (predicting $X$ from $Y$) is symmetric and follows by swapping roles.
@@ 4 Bridged Clustering Algorithm
Input: Input-only set $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{|\mathcal{X}|}\}$; Output-only set $\mathcal{Y}=\{y_{1},y_{2},\ldots,y_{|\mathcal{Y}|}\}$; sparse supervised set $\mathcal{S}=\{(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{k}$; Test-set $\mathcal{X_{\text{test}}}=\{\hat{x}_{1},\hat{x}_{2},\ldots,\hat{x}_{|\mathcal{X_{\text{test}}}|}\}$
@@
1: Apply a clustering algorithm to $\mathcal{X}$ (respectively, $\mathcal{Y}$) to obtain cluster assignments $\mathcal{C}_{\mathcal{X}}$ (respectively, $\mathcal{C}_{\mathcal{Y}}$).
2: Using the supervised set $\mathcal{S}$, learn a mapping $\hat{A}_{x\to y}$ between clusters in $\mathcal{X}$ and clusters in $\mathcal{Y}$.
3:for each sample $x_{i}$ in $\mathcal{X_{\text{test}}}$ do
4:  Assign $x_{i}$ to a cluster $c_{x}=\mathcal{C}_{\mathcal{X}}(x_{i})$.
5:  Find corresponding cluster in $\mathcal{Y}$: $c_{y}=\hat{A}_{x\to y}(c_{x})$.
6:  Predict $\hat{y}_{i}$ as the centroid of cluster $c_{y}$ in $\mathcal{Y}$.
7:end for
8:return Predicted outputs $\hat{\mathcal{Y}}$ for test inputs $\mathcal{X_{\text{test}}}$
@@
We first run a $k$-clustering algorithm to partition the dataset into $k$ groups separately on $\mathcal{X}$ and $\mathcal{Y}$ using $C$ centers:
$\hat{c}_{X}:\mathbb{R}^{d}\!\to[C],\hat{c}_{Y}:\mathbb{R}^{d^{\prime}}\!\to[C].$
Because cluster labels are only defined up to a permutation of indices, we evaluate their quality via the permutation-invariant mis-clustering rates:
$\varepsilon_{X}:=\min_{\sigma_{X}:[C]\!\to\![C]}\Pr\bigl[\sigma_{X}(\hat{c}_{X}(X))\neq T\bigr]$
$\varepsilon_{Y}:=\min_{\sigma_{Y}:[C]\!\to\![C]}\Pr\bigl[\sigma_{Y}(\hat{c}_{Y}(Y))\neq T\bigr].$
Let $\sigma_{X},\sigma_{Y}$ denote permutations that attain these minima, and let $\sigma_{X}^{-1},\sigma_{Y}^{-1}$ denote their inverses.